# Fantastic Semantics and Where to Find Them: Investigating Which Layers of Generative LLMs Reflect Lexical Semantics

### v1

@@ Abstract
Large language models have achieved remarkable success in general language understanding tasks. However, as a family of generative methods with the objective of next token prediction, the semantic evolution with the depth of these models are not fully explored, unlike their predecessors, such as BERT-like architectures. In this paper, we specifically investigate the bottom-up evolution of lexical semantics for a popular LLM, namely Llama2, by probing its hidden states at the end of each layer using a contextualized word identification task. Our experiments show that the representations in lower layers encode lexical semantics, while the higher layers, with weaker semantic induction, are responsible for prediction. This is in contrast to models with discriminative objectives, such as mask language modeling, where the higher layers obtain better lexical semantics. The conclusion is further supported by the monotonic increase in performance via the hidden states for the last meaningless symbols, such as punctuation, in the prompting strategy.
@@ 1 Introduction
Figure 1: Key Differences between BERT and Llama2 Language Models. Blue and red lines indicate the information flows of understanding and predicting.
Given these observations, we hypothesize that GPT-like LLMs encode lexical semantics in lower layers while making predictions, potentially leading to the forgetting of information related to current tokens in higher layers. This hierarchical behavior suggests a dynamic interaction between understanding and prediction in generative LLMs, as indicated by the view of information flow in recent studies Wang et al. (2023a); Voita et al. (2019).
@@ 3.2 Settings and Models
For a given word $w$ within context $c$, Llama2 extracts hidden states $h_{i}\in\mathbb{R}^{D}$ across each of its 32 layers, where $D$ is 4096 in Llama2. The cosine similarity of $w$ in paired contexts ($c^{a}$, $c^{b}$) is calculated as $s^{ab}_{w}$. Subsequently, sentence pairs are classified as true if $s^{ab}_{w}$ exceeds a threshold $\gamma$, and false if it falls below $\gamma$. The optimal $\gamma$ is determined through development dataset, with distinct values potentially assigned for each layer to accommodate varying similarity ranges. The optimal values of $\gamma$ are listed in Appendix A.1. To address potential anisotropy in the embedding space, we employ standardization across samples following prior research Ethayarajh (2019).
We employ different input variants for Llama2. The base setting uses the original context $c$ with lexical representations $h_{i}$ at the target position. Since $w$ cannot access the context behind it in this setting, we repeat the original context and obtain $h_{i}$ in the second context, ensuring all information is left of $w$. This configuration is referred to as repeat. Another setting is inspired by the prompting strategy proposed in the paper Jiang et al. (2023). Here, we modify the context $c$ as: The $w$ in this sentence: $c$ means in one word :. Then, we calculate the representation from the position of the last token, i.e., the final colon :, as $h_{i}$ and we denote this as prompt. An example is provided in Table 1.
@@
base
the bank of the river
repeat
the bank of the river the bank of the river
prompt
The bank in this sentence:“the bank of the river” means in one word :
Table 1: An example to show different input formats in three settings. Bold token positions are used as hidden states $h_{i}$ of target words.
@@ 4 Results and Analysis
Llama2_base†(6) 60.9 62.6 62.1
Llama2_base (12) 62.9 63.9 62.5
Llama2_repeat†(8) 64.6 66.5 63.6
Llama2_repeat (7) 68.4 70.3 65.5
Llama2_prompt†(23) 68.9 68.9 69.5
Llama2_prompt (20) 69.2 69.2 69.5
Table 2: Overall accuracy (%) on the WiC test set. †indicates methods without anisotropy removal. The numbers in brackets after the model name indicate the number of layers for achieving the best performance.
Table 2 presents the overall performance. Llama2, as a generative model, achieves comparable results to bidirectional and non-regressive BERT models, outperforming non-transformer models like Elmo. This suggests that LLMs have the potential for word-level understanding, even though it is not explicitly trained for this capability. As expected, the prompting strategy achieves the highest accuracy among all the Llama2 variants. This approach incorporates downstream tasks into the generative process during LLM training and has proven to be popular and effective in addressing both intermediate and high-level tasks in the LLM era Zhao et al. (2023). However, prompting relies on the choice of prompts and may not directly reveal the model’s internal understanding. On the other hand, our repeat strategy demonstrates comparable performance to prompting and significantly outperforms the base version (with a 5.5 advantage gap). This simple yet effective transformation strikes a balance between information accessibility and prompting robustness.
In terms of parts-of-speech, nouns generally exhibit higher accuracy than verbs, as evidenced by a 4.8 advantage gap in Llama2_repeat. This observation aligns with previous studies that have concluded that verbs are more challenging to disambiguate Barba et al. (2021).
@@
Figure 2: Layer-wise accuracy for different settings and models. Star shows the best value.
@@
Figure 2 illustrates the layer-wise dynamics in two settings for Llama2 and also BERT_large. We observe non-monotonic trends for Llama2 across layers: both the base and repeat initially increase in lower layers before decreasing in higher layers. Consequently, optimal performance is achieved at lower layers when utilizing the hidden states of the target word as the default choice. This suggests that lower layers in LLMs encode lexical semantics, offering both a practical insight and a pathway for interpreting LLMs. Moreover, the trend contrasts with bidirectional BERT_large model, which obtains the best performance in higher layers. This highlights a difference between these two architectures: BERT concentrates on its current word across the layers while Llama2 aims for next token prediction.
@@ 5 Conclusion
This study investigates how Llama2’s layer-wise representations encode lexical semantics using the WiC dataset. Our experiments reveal that optimal performance is achieved at lower layers for generative tasks, while predictive accuracy improves in higher layers. This suggests that Llama2 prioritizes understanding before prediction as information flows from lower to higher layers. These findings offer practical guidance on extracting representations for lexical semantics tasks in engineering applications and shed light on the interpretability of LLMs due to their impressive performance.

### v2

@@ Abstract
Large language models have achieved remarkable success in general language understanding tasks. However, as a family of generative methods with the objective of next token prediction, the semantic evolution with the depth of these models are not fully explored, unlike their predecessors, such as BERT-like architectures. In this paper, we specifically investigate the bottom-up evolution of lexical semantics for a popular LLM, namely Llama2, by probing its hidden states at the end of each layer using a contextualized word identification task. Our experiments show that the representations in lower layers encode lexical semantics, while the higher layers, with weaker semantic induction, are responsible for prediction. This is in contrast to models with discriminative objectives, such as mask language modeling, where the higher layers obtain better lexical semantics. The conclusion is further supported by the monotonic increase in performance via the hidden states for the last meaningless symbols, such as punctuation, in the prompting strategy. Our codes are available at https://github.com/RyanLiut/LLM_LexSem.
@@ 1 Introduction
Figure 1: Key Differences between BERT and Llama2 Language Models. Blue and red lines indicate the information flows of understanding and predicting. The "understanding" refers to capture the lexical semantics by leveraging context. The blue line from the context to the current word indicates the flow of understanding.
@@
In addition, the extraction of contextual word representations from generative Large Language Models (LLMs) often proves to be more beneficial than those derived from models akin to BERT in certain scenarios. Firstly, as generative LLMs are increasingly recognized for their robustness and prevalence as foundational models Zhao et al. (2023), we only obtain “autoregressive embeddings” without compromising their generative capabilities. Secondly, these LLMs are consistently trained with billions of parameters across expansive web-scale datasets, thereby exhibiting a significantly greater potential and anticipated superior performance compared to their much lighter counterparts, such as BERT-like models.
Given these observations, we hypothesize that GPT-like LLMs encode lexical semantics in lower layers while making predictions, potentially leading to the forgetting of information related to current tokens in higher layers. [^1] We refer to lower layers as those closer to the inputs, while higher layers as closer to the outputs. This hierarchical behavior suggests a dynamic interaction between understanding and prediction in generative LLMs, as indicated by the view of information flow in recent studies Wang et al. (2023a); Voita et al. (2019).
@@ 3.2 Settings and Models
For a given word $w$ [^3] We average the hidden states for tokens within the word as the final word representation. within context $c$, Llama2 extracts hidden states $h_{i}\in\mathbb{R}^{D}$ across each of its 32 layers, where $D$ is 4096 in Llama2. The cosine similarity of $w$ in paired contexts ($c^{a}$, $c^{b}$) is calculated as $s^{ab}_{w}$. Subsequently, sentence pairs are classified as true if $s^{ab}_{w}$ exceeds a threshold $\gamma$, and false if it falls below $\gamma$. The optimal $\gamma$ is determined through development dataset, with distinct values potentially assigned for each layer to accommodate varying similarity ranges. The optimal values of $\gamma$ are listed in Appendix A.1. To address potential anisotropy in the embedding space, we employ standardization across samples following prior research Ethayarajh (2019).
We employ different input variants for Llama2. The base setting uses the original context $c$ with lexical representations $h_{i}$ at the target position. Since $w$ cannot access the context behind it in this setting, we repeat the original context and obtain $h_{i}$ in the second context, ensuring all information is left of $w$. This configuration is referred to as repeat. We also explore the word before the target one to valid the predictive ability in higher layers, which is denoted as repeat_prev. Another setting is inspired by the prompting strategy proposed in the paper Jiang et al. (2023). Here, we modify the context $c$ as: The $w$ in this sentence: $c$ means in one word :. Then, we calculate the representation from the position of the last token, i.e., the final colon :, as $h_{i}$ and we denote this as prompt. An example is provided in Table 1.
@@
base the bank of the river
repeat the bank of the river the bank of the river
repeat_prev the bank of the river the bank of the river
prompt In this sentence “the bank of the river”, “bank” means in one word :
Table 1: An example to show different input formats in three settings. Bold token positions are used as hidden states $h_{i}$ of target words. We highlight that the bold final colon in the prompt setting is used to extract $h_{i}$.
@@ 4 Results and Analysis
Llama2_base†(6) 60.9 63.7 58.3
Llama2_base (11) 63.6 66.8 58.7
Llama2_repeat†(9) 64.5 66.4 63.4
Llama2_repeat (8) 68.1 72.7 65.6
Llama2_prompt†(28) 71.1 68.9 72.9
Llama2_prompt (21) 72.7 74.5 72.1
Table 2: Overall accuracy (%) on the WiC test set. †indicates methods without anisotropy removal. The numbers in brackets after the model name indicate the number of layers for achieving the best performance. The index begins at 0, representing the input embedding layer and increases as the model goes deeper. This applies similarly to the remaining indices in the figures.
Table 2 presents the overall performance. Llama2, as a generative model, achieves comparable results to bidirectional and non-regressive BERT models, outperforming non-transformer models like Elmo. This suggests that LLMs have the potential for word-level understanding, even though it is not explicitly trained for this capability. As expected, the prompting strategy achieves the highest accuracy among all the Llama2 variants. This approach incorporates downstream tasks into the generative process during LLM training and has proven to be popular and effective in addressing both intermediate and high-level tasks in the LLM era Zhao et al. (2023). However, prompting relies on the choice of prompts and may not directly reveal the model’s internal understanding. On the other hand, our repeat strategy demonstrates comparable performance to prompting and significantly outperforms the base version (with a 4.5 advantage gap). This simple yet effective transformation strikes a balance between information accessibility and prompting robustness.
In terms of parts-of-speech, nouns generally exhibit higher accuracy than verbs, as evidenced by a 7.1 advantage gap in Llama2_repeat. We also observe that in the base setting, verbs exhibit significantly lower accuracy, with decreases of 8.1 points compared to nouns. This decline is attributed to the fact that disambiguating verbs requires more context, which is often lacking in real data preceding the verbs. For instance, target verbs positioned at the beginning of sentences, where there is no prior context to aid in disambiguation, account for 19.2% of cases, in contrast to 14.3% for nouns. These observations align with previous studies that have concluded that verbs are more challenging to disambiguate Barba et al. (2021).
@@
Figure 2: Layer-wise accuracy for different settings and models (Llama2 and BERT_large). Star shows the best value.
@@
Figure 2 illustrates the layer-wise dynamics in two settings for Llama2 and also BERT_large. We observe non-monotonic trends for Llama2 across layers: both the base and repeat initially increase in lower layers before decreasing in higher layers. Consequently, optimal performance is achieved at lower layers when utilizing the hidden states of the target word as the default choice. This suggests that lower layers in LLMs encode lexical semantics, offering both a practical insight and a pathway for interpreting LLMs. Notice that the performance in the highest layer for LLMs does not lose to the worst (i.e., 50%), indicating it has still remained word meaning in some extent. Moreover, the trend contrasts with bidirectional BERT_large model, which obtains the best performance in higher layers. This highlights a difference between these two architectures: BERT concentrates on its current word across the layers while Llama2 aims for next token prediction.
@@ 5 Conclusion
This study investigates how Llama2’s layer-wise representations encode lexical semantics using the WiC dataset. Our experiments reveal that optimal performance is achieved at lower layers for generative tasks, while predictive accuracy improves in higher layers. This suggests that Llama2 prioritizes understanding before prediction as information flows from lower to higher layers. These findings offer practical guidance on extracting representations for lexical semantics tasks in engineering applications. For example, we would opt to utilize representations from the lower layers for lexical-related tasks, such as POS tagging and word sense disambiguation. Conversely, those from the higher layers could be employed for prediction-related or generative tasks, including text summarization and dialogue generation. Furthermore, it also sheds light on the interpretability of LLMs from a top-down perspective.
@@ 7 Ethics Statement
8 Broader Impact Statement
Understanding the linguistic knowledge that LLMs have acquired is fundamental for the practical application of generative AI in the real world. This understanding not only enhances the interpretability of these black-box “Goliaths,” but also improves the robustness, reliability, and safety of the models. Words carry significant linguistic meaning, while the counterpart tokens serve as the smallest computational units for transformers. We believe that exploring the lexical semantics within LLMs is a foundational step in bridging the gap between computational modeling and linguistics, thereby highlighting the benefits of combining both fields.
9 Acknowledgements
The authors thank the anonymous reviewers for their valuable comments and constructive feedback on the manuscript. We also thank members of THUNLP Lab for their valuable discussions. This work is supported by the 2018 National Major Program of Philosophy and Social Science Fund “Analyses and Researches of Classic Texts of Classical Literature Based on Big Data Technology” (18ZDA238) and Research on the Long-Term Goals and Development Plan for National Language and Script Work by 2035 (ZDA145-6).